### v4

# Probabilistic Analysis of Copyright Disputes and Generative AI Safety

Abstract.
This paper presents a probabilistic approach to analyzing copyright infringement disputes. Under this approach, evidentiary principles shaped by case law are formalized in probabilistic terms, allowing for a mathematical examination of issues arising in such disputes. The usefulness of this approach is showcased through its application to the “inverse ratio rule”—a controversial legal doctrine adopted by some courts. Although this rule has faced significant criticism, a formal proof demonstrates its validity, provided it is properly defined. Furthermore, the paper employs the probabilistic approach to study the copyright safety of generative AI. Specifically, the Near Access-Free (NAF) condition, previously proposed as a strategy for mitigating the heightened copyright infringement risks of generative AI, is evaluated. The analysis reveals that, while the NAF condition mitigates some infringement risks, its justifiability and efficacy are questionable in certain contexts. These findings illustrate how taking a probabilistic perspective can enhance our understanding of copyright jurisprudence and its interaction with generative AI technology.
copyright: noneccs: Social and professional topics Copyrightsccs: Computing methodologies Probabilistic reasoning
1. Introduction
Copyright law grants creators exclusive rights to their original works across diverse categories of creative expression, such as literature, music, and art, with the aim of incentivizing creation and enriching society (Gorman, 2006). The rise of new technologies has continuously challenged conventional copyright principles, and in response, copyright jurisprudence has evolved over time, becoming increasingly complex (Joyce et al., 2020). Despite the ever-growing need to understand how copyright law interacts with emerging technologies—generative AI being a prominent recent example (Samuelson, 2023; Lemley, 2024)—the theoretical framework remains underdeveloped, leading to significant legal uncertainty (Depoorter, 2009).
This article employs a probabilistic approach to analyze copyright infringement disputes, seeking to bring clarity to its complexities. The analysis is carried out by distilling and formalizing the fundamental principles governing copyright disputes as established through case law in the United States, particularly focusing on the requirements a copyright owner must meet to make a prima facie case of copyright infringement in court. This body of jurisprudence encompasses not only the substantive elements of copyright law, but also the evidentiary principles that guide how evidence is presented and assessed (Loren and Reese, 2019).
The theoretical foundation of this work lies in the long-recognized link between probabilistic reasoning and judicial proof (Eggleston, 1989; Cullison, 1969; Broun and Kelly, 1970; Finkelstein and Fairley, 1970; Cohen, 1977; Eggleston, 1978; McCauliff, 1982). Specifically, the legal fact-finding process can be understood to involve assessment of the probabilities of allegations being true or false based on the available evidence (Stein, 2005a; Cheng, 2013). Building on this link, core evidentiary principles in copyright jurisprudence can be formalized in probabilistic terms, allowing for a mathematical analysis of a controversial legal doctrine, the “inverse ratio rule.” This rule governs the relationship between two critical pieces of evidence in copyright infringement cases—access and similarity to the original work. Although the validity of this rule has been the subject of considerable debate (Aronoff, 2007; Balganesh and Menell, 2022), a formal proof is established here, demonstrating that it holds under natural assumptions when its content and scope are carefully delineated.
Furthermore, this approach facilitates the analysis of copyright infringement risks introduced by technological advancements and the identification of effective mitigation strategies. Specifically, the probabilistic approach is utilized to evaluate the efficacy of the Near Access-Free (NAF) condition (Vyas et al., 2023), which has been proposed to mitigate the risk of copyright infringement posed by generative AI. This paper identifies extensive access to copyrighted materials by generative models as a significant source of heightened infringement risks. While the NAF condition addresses this issue to some extent, the analysis reveals that it does so partly on ethically questionable grounds—namely, by reducing transparency regarding the materials to which generative AI had access during training—and that its effectiveness in mitigating infringement risks remains limited under certain situations.
2. Related Work
Existing research has investigated the relationship between copyright law and probabilistic thinking, specifically concerning how the probability of creation relates to concepts of creativity and originality within the context of copyright law (Byron, 2006; Khong, 2007; Vermont, 2012; Hacohen and Elkin-Koren, 2024; Chiba-Okabe and Su, 2024). However, there has been little discussion on the application of probabilistic approaches in analyzing the evidentiary rules of copyright infringement disputes. While a previous study has alluded to the connection between the probabilistic nature of legal fact-finding and the inverse ratio rule (Balganesh and Menell, 2022), no formal probabilistic analysis of the rule has been undertaken, to the best of the author’s knowledge. This gap is addressed by employing a probabilistic approach to formalize the evidentiary principles governing copyright disputes and mathematically analyze the inverse ratio rule.
In response to rapid advancements in generative AI technology, a substantial body of literature has emerged, addressing potential copyright risks and defenses available to developers (Sag, 2023; Henderson et al., 2023; Lee et al., 2024). These works discuss various legal challenges, including the potential for copyright infringement arising from AI-generated content, with some studies analyzing the Near Access-Free (NAF) condition as a mitigation strategy (Lee et al., 2024; Elkin-Koren et al., 2024). This paper contributes to the discourse by providing a probabilistic analysis of legal issues surrounding the NAF condition, particularly through the lens of inference based on evidence in copyright infringement disputes.
3. Analysis of Copyright Disputes
3.1. Essential Evidentiary Principles in Copyright Disputes
To prevail in a copyright infringement dispute, the copyright owner, as the plaintiff, must demonstrate two elements under the traditional Arnstein test: “(a) that defendant copied from plaintiff’s copyrighted work and (b) that the copying (assuming it to be proved) went so far as to constitute improper appropriation.” [^1] Arnstein v. Porter, 154 F.2d 464, 468 (2d Cir. 1946). These two requirements are often referred to simply as “factual copying” and “improper appropriation,” respectively (Balganesh and Menell, 2022). Although several defenses, including fair use, may be available to the defendant in some cases even if these elements are established (Loren and Reese, 2019), these defenses are not considered in this paper to maintain focus on the essential structure of copyright infringement cases.
Direct evidence of factual copying includes admission by the defendant or an eyewitness account of the replication or appropriative behavior, but such evidence is hard to obtain. [^2] See, e.g., Smith v. Jackson, 84 F.3d 1213, 1218 (9th Cir. 1996), stating “direct evidence of copying is not available in most cases.” Therefore, factual copying is almost invariably proved by circumstantial evidence (Gorman, 2006; Balganesh and Menell, 2022), and the evidence generally accepted by courts consists of that pertaining to access and similarity. Specifically, copying is proved based on (i) evidence showing that the defendant had access to the original work and (ii) the similarity between the original work and the defendant’s work (Balganesh and Menell, 2022; Loren and Reese, 2019). Courts generally hold that proof of access combined with a certain level of similarity, that is probative of copying, is sufficient to establish factual copying (Latman, 1990; Loren and Reese, 2019). Although both access and similarity support the inference of factual copying, simply having proof of access does not suffice to establish factual copying when there is no similarity between the two works. [^3] Arnstein, 154 F.2d at 468. Proof of access pertains to the antecedent of factual copying, in the sense that the defendant having had access to the original work is a necessary condition for engaging in copying (Balganesh and Menell, 2022; Latman, 1990). In contrast, similarity emerges as a result of copying (Balganesh and Menell, 2022; Latman, 1990), and when the similarity between the two works is so pronounced as to be deemed “striking,” this alone suffices to establish that copying has occurred. [^4] Id. at 468-69. In such cases, it can be inferred that the defendant had access to the original work based on the high degree of similarity. [^5] Ty, Inc. v. GMA Accessories, Inc., 132 F.3d 1167, 1170 (7th Cir. 1997), stating “a similarity that is so close as to be highly unlikely to have been an accident of independent creation is evidence of access.”
The second requirement, improper appropriation, is established by demonstrating that the extent of copying reaches the threshold at which the similarity between the original work and the defendant’s work can be deemed “substantial” in degree. Determination of this requirement involves a normative judgment, taking into account both quantitative and qualitative aspects of copying (Gorman, 2006; Balganesh, 2012). Thus, similarity serves a dual role: it supports the finding of factual copying and acts as an independent criterion for establishing improper appropriation (Latman, 1990; Balganesh and Menell, 2022), though the sense of similarity differs in each context, even if there may be some overlap (Loren and Reese, 2019).
3.2. Formalization of Evidentiary Principles
Legal fact-finding is a reasoning process that involves assessing and weighing available evidence to reach a decision amid uncertainty. This process may be appropriately characterized by probabilistic reasoning, wherein adjudicators employ common sense and logic, drawing on knowledge of the frequencies of occurrences of certain events and patterns derived from experience to evaluate the likelihood of the particular circumstances in question (Stein, 2005b). Although the absolute truth typically cannot be attained, the final decision in most civil cases is made based on the standard of “preponderance of the evidence.” This standard requires the party bearing the burden of proof to demonstrate that their claim is more likely true than not. From a probabilistic perspective, it is often understood as requiring a demonstration of greater than 50% likelihood that the claim is true (McCauliff, 1982; Stein, 2005a), particularly in binary situations where the question is whether an assertion is true or false (Cheng, 2013). In criminal cases, the relevant standard is “proof beyond a reasonable doubt,” which is commonly understood to require a significantly higher probability (McCauliff, 1982). Probabilistic approaches to fact-finding have been extensively explored in the literature over many years (Eggleston, 1989; Cullison, 1969; Broun and Kelly, 1970; Finkelstein and Fairley, 1970; Cohen, 1977; Eggleston, 1978; McCauliff, 1982) and, despite considerable criticisms (Tribe, 1971; Allen and Stein, 2013), have remained influential, even if primarily at a conceptual level.
Building upon this foundation, the essential structure of copyright disputes can be formalized by representing adjudicators’ degree of belief in terms of subjective probabilities, utilizing a discrete probability space $(\Omega,\mathcal{F},P)$ and random variables defined on this space. The condition for copyright infringement to exist can be expressed using random variables as $Copy=1\wedge Substantial=1$. Here, $Copy$ is a binary random variable representing the presence ($Copy=1$) or non-presence ($Copy=0$) of factual copying of the plaintiff’s copyrighted work during the creation of the defendant’s work. Similarly, $Substantial$ represents the presence or non-presence of substantial similarity between the two works. Typically, adjudicators can determine whether substantial similarity exists, or the truth value of $Substantial$, with certainty by directly comparing the original work to the defendant’s work presented at trial. In contrast, the absolute truth about whether the defendant engaged in factual copying remains unknown during the trial, necessitating a probabilistic approach to account for this uncertainty. The plaintiff must present sufficient evidence for the adjudicators to conclude that the probability of $Copy=1$ exceeds a certain threshold, $\lambda$, which could be 0.5, aligning with the preponderance of evidence standard.
The case law establishing that probative similarity combined with access typically suffices to prove copying can be represented in probabilistic terms as follows:
$\begin{gathered}P(Copy=1|Access=1,\,Probative=1)=\eta\\ \textrm{such that}\,\,\eta>\lambda\end{gathered}$
where $Access$ and $Probative$ are binary random variables representing whether access by the defendant to the plaintiff’s original work was present and whether the original and defendant’s works are probatively similar, respectively. Suppose the adjudicators learn from the evidence presented during the trial that the defendant indeed had access to the original work ($Access=1$) and that the defendant’s work is probatively similar to the original work ($Probative=1$). This results in the adjudicators’ degree of belief that copying occurred ($Copy=1$) surpassing the threshold $\lambda$ (since $\eta>\lambda$), thereby the standard of proof being met, as desired.
This approach can accommodate the logical relationship between factual copying and access, specifically that the defendant must have had access to the original work in order to copy from it. This relationship can be represented as:
$P(Copy=1|Access=0)=0.$
As a result of imposing this restriction on $Copy$, if it is known with certainty that the defendant did not have access to the original work, the degree of belief that the defendant copied the original work must be zero, as expected.
A further step can be taken to represent the evidentiary rule that striking similarity is typically sufficient to establish copying as follows:
$\begin{gathered}P(Copy=1|Access=1,\,Striking=1)=\rho\\ P(Access=1|Striking=1)=\sigma\\ \textrm{such that}\,\,\rho\sigma>\lambda\end{gathered}$
where $Striking$ is a binary random variable corresponding to the presence or non-presence of striking similarity. Assuming that the adjudicators learn during the trial that $Striking=1$ with certainty, taking into account the logical relationship between access and copying, the degree of belief about copying can be computed as:
$\begin{split}&P\left(Copy=1|Striking=1\right)\\ &=P\left(Copy=1|Striking=1,\,Access=1\right)P\left(Access=1|Striking=1\right)\\ &\,\,+P\left(Copy=1|Striking=1,\,Access=0\right)P\left(Access=0|Striking=1\right)\\ &=P\left(Copy=1|Striking=1,\,Access=1\right)P\left(Access=1|Striking=1\right)\\ &\,\,+0\cdot P\left(Access=0|Striking=1\right)\\ &=\rho\sigma>\lambda.\end{split}$
Therefore, the degree of belief of the adjudicators reaches the threshold corresponding to the applicable standard of proof, allowing them to conclude that copying has occurred.
3.3. The Inverse Ratio Rule
The so-called “inverse ratio rule,” adopted by some courts, dictates a specific relationship between proof of access and similarity: stronger proof of access lowers the level of similarity required to establish copyright infringement, while a higher level of similarity reduces the burden of proof regarding access (Aronoff, 2007; Balganesh and Menell, 2022). Although the rule has been applied for nearly a century, serious objections have been raised to its application (Aronoff, 2007). The controversial status of the inverse ratio rule gained renewed attention with its recent abrogation by the Ninth Circuit Court, [^6] Skidmore v. Led Zeppelin, 952 F.3d 1051, 1066-69 (9th Cir. 2020) (en banc). while it remains in use in some other courts (Balganesh and Menell, 2022).
The controversy surrounding the inverse ratio rule partly arises from a likely inadvertent divergence in the understanding of its scope and content. Some courts have applied the inverse ratio rule in a way that strong evidence of access lowers the level of similarity required to establish the improper appropriation prong necessary for proving infringement (Balganesh and Menell, 2022; Aronoff, 2007). However, there appears to be little justification for the inverse ratio rule operating in this manner (Balganesh and Menell, 2022). Another, more promising formulation of the inverse ratio rule pertains to the factual copying prong, where the rule adjusts the degree of similarity or the strength of evidence of access required to establish factual copying. Specifically, when strong evidence of access is present, a lower degree of similarity is sufficient to prove factual copying, whereas weak evidence of access necessitates a higher level of similarity. Conversely, a high degree of similarity lessens the evidence of access required, and vice versa. Proponents argue that the inverse ratio rule of this form follows from general principles of evidential inference (Balganesh and Menell, 2022). However, critics contend that even the inverse ration rule formulated in this way is flawed, as access is merely a necessary condition for copying, and better evidence of access does not increase the probability of copying (Aronoff, 2007). This formulation of the inverse ratio rule was also the one rejected by the Ninth Circuit Court. [^7] Id. at 1066-68.
The analysis here precisely describes the content of the inverse ratio rule as it pertains to the factual copying element in formal language and proves that the rule, defined in this way, must hold. The proof relies solely on assumptions that are naturally deduced from well-established jurisprudence. In particular, we consider a set of random variables that encapsulates the evidentiary principles discussed in the preceding section in a more generalized form (Definition 3.1).
Definition 3.1.
$C,A\in\{0,1\}$, $S\in\{1,2,\cdots,n\}$, and $EA\in\{1,2,\cdots,m\}$ are random variables that satisfy:
(1) $P(C=1|A=0)=0$
(2) $P(C=1|A=1,\,S=i)=\alpha_{i},\,\forall i$
(3) $P(A=1|S=i,\,EA=j)=\beta_{i,j},\,\forall(i,j)$
(4) $EA\perp\!\!\!\perp\{C,S\}|A$
where $\alpha_{i+1}\geq\alpha_{i}$ for all $i$, $\beta_{i,j+1}\geq\beta_{i,j}$ for all $j$ with $i$ fixed, and $\beta_{i+1,j}\geq\beta_{i,j}$ for all $i$ with $j$ fixed.
$C$ and $A$ are abbreviations of $Copy$ and $Access$, respectively, from earlier discussions, which enjoy a logical relationship (Definition 3.1 (1)). Possible values of $S$ represent varying degrees of similarity between the original and defendant’s works in the probative sense. As reflected in the well-established evidentiary rule that copying can be established by demonstrating both access and probative similarity, similarity serves as significant evidence in proving factual copying (Definition 3.1 (2)). It is reasonable, at least in principle (assuming no other specific circumstances suggest otherwise), to think that there is a monotonic relationship among the inferential values of different degrees of similarity; that is, a higher degree of similarity more strongly infers factual copying, so that $\alpha_{i+1}\geq\alpha_{i}$ holds. Similarity also has evidentiary value in establishing access to the original work (Definition 3.1 (4)), as demonstrated by the evidentiary rule regarding striking similarity. Even critics of the inverse ratio rule acknowledge the evidentiary value of similarity for establishing access, though they suggest that the strength of the inference may be weaker than assumed by others (Aronoff, 2007). It is natural to posit that such evidentiary value for access increases monotonically with similarity in principle, leading to the assumption that $\beta_{i+1,j}\geq\beta_{i,j}$ for all $i$.
Variable $EA$ is introduced to capture the interaction between levels of similarity and evidence of access (other than similarity itself). The values it can take correspond to varying degrees of the strength of evidence of access presented in court (Definition 3.1 (3)). For example, strong evidence of access might include testimony that the original work was directly handed to the defendant or the fact that the work was publicly accessible and widely known. Conversely, the fact that the original work was kept private, with little chance of the defendant encountering it, would imply a low likelihood of access. Such evidence is assumed to have direct impact only on the inference of access and is independent of other variables given $A$ (Definition 3.1 (4)). The differing strengths of evidence naturally create an ordered variation in the inferential value of access, captured by the relationship $\beta_{i,j+1}\geq\beta_{i,j}$ for all $j$.
During court proceedings, adjudicators learn both the level of similarity between the works and the strength of evidence of access presented, corresponding to a pair of values $(i,j)$. Then, the degree of belief of the adjudicators about the presence of factual copying, $C=1$, are formed given the knowledge of $S=i$ and $EA=j$.
Under these assumptions, the inverse ratio rule as it pertains to factual copying can now be precisely articulated in formal language and its correctness proved, as summarized in Proposition 3.2 (see Appendix C for proofs).
Proposition 3.2.
The following inverse ratio rule holds for any $\lambda\in(0,1)$:
  • If $j\geq j^{\prime}$, $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$
  • If $i\geq i^{\prime}$, $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$
where the random variables are defined as in Definition 3.1.
The fact that access constitutes a necessary condition for copying does not imply that the strength of the evidence of access cannot affect the inference of copying, as critics of the inverse ratio rule argue. Stronger evidence of access can increase the likelihood of copying because the likelihood of copying is the product of two factors: (i) the likelihood of copying given the presence of access and the level of similarity, and (ii) the likelihood of access given the level of similarity and other evidence of access. Thus, stronger evidence of access leads to an increased likelihood of copying by elevating the value of the second factor in the product. Although Proposition 3.2 shows the inverse ratio rule only with weak inequalities, which can be proved with minimal assumptions, the interplay between these two factors allows the inverse ratio rule to hold in an even stronger sense, with strict inequalities, in certain specific yet highly plausible circumstances (see Appendix A for further details).
Furthermore, although the conclusion that the inverse ratio rule holds true is accepted, the analysis indicates that the reasoning posited by its proponents does not fully capture the intricacies involved. The argument previously put forward to support the inverse ratio rule contends that two pieces of relatively weak evidence—pertaining to access and similarity, respectively—possess strong inferential value when combined (Balganesh and Menell, 2022). However, the reasoning of the results concerning the inverse ratio rule demonstrates that the level of similarity and the strength of evidence on access interact also in a complementary manner, rather than merely cumulatively, due to their multiplicative effect on the probability of copying. Specifically, a higher degree of similarity increases factor (i), and even when combined with a relatively lower factor (ii) resulting from weaker evidence of access, the product of these probabilities can still surpass the necessary threshold for establishing factual copying, and vice versa. Since the product of two probabilities is always smaller than either individual factor, combining small factors—stemming from weak evidence in both similarity and access—cannot yield a stronger inference on its own. This demonstrates that the cumulative inference logic is insufficient to fully explain the interaction between evidence of access and similarity.
One might notice the product of the two probability terms is reminiscent of the so-called conjunction problem that has been a point of debate between legal scholars who support probabilistic views of legal fact-finding and those who are skeptical about these views. However, as elaborated in Appendix B, the situation at hand is different from typical scenarios that are considered in the context of the conjunction problem, and discussions surrounding the problem does not affect the analysis of the inverse ratio rule.
4. Copyright Safety for Generative AI
4.1. Evidentiary Considerations in Copyright Disputes Involving AI-Generated Works
When generative AI is used to create content, it is typically offered as a service by the AI developer and used by a user. In such cases, both the user and the AI developer may be considered direct or indirect infringers if the generated output is substantially similar to an original work, depending on the specific circumstances. Copyright infringement may occur even without the user or developer intending to replicate the copyrighted material or to create something similar (Lee et al., 2024).
Due to the uninterpretability of generative models, it can be challenging to provide direct evidence of factual copying for AI-generated content. In most cases, access and similarity will likely serve as the primary evidence of infringement, much like in traditional copyright disputes (Lee et al., 2024). While the creative processes of generative models differ markedly from human creativity, it appears that similar evidentiary principles are likely to apply even when generative AI is involved in the creative processes. At a minimum, it is generally recognized that copying, even for generative AI outputs, requires some form of access to the original work, which may be demonstrated by its inclusion in the AI model’s training data (Zirpoli, 2023; Lee et al., 2024; Elkin-Koren et al., 2024). Moreover, studies have shown that exposure to certain types of content during training leads to a higher likelihood of these models reproducing or closely imitating that content (Carlini et al., 2021, 2023; Hacohen et al., 2024), suggesting that inclusion of specific data correlates with a greater probability of producing similar outputs through copying. This echoes the evidentiary framework applicable to traditional copyright cases, where greater similarity between works strengthens the inference of access and copying.
A distinctive feature of generative models, setting them apart from human artists and traditional creative tools, is their capacity to be trained on vast and diverse datasets. Training datasets are often compiled by scraping content from the Internet, including copyrighted material (Lee et al., 2024). Furthermore, direct inclusion of copyrighted material in the training data is not even necessary for latent diffusion models to produce images similar to copyrighted works, as these models can reproduce content from other images that retain similar latent information (Lu et al., 2024), which exacerbates the problem. These characteristics suggest a high likelihood that generative models have what can be construed as “access” to copyrighted works. As a result, knowing that generative AI was used to create the defendant’s work would likely lead to a stronger belief in the presence of access. For example, consider a scenario where the original work was posted online, and the defendant’s work was created by an AI model trained on data scraped from the Internet. Even in the Internet age, simply posting a work online does not automatically establish access by the defendant. Courts require more than just online availability; factors like the work’s visibility or significant dissemination are needed to strengthen the inference of access. [^8] See generally Cooley v. Target Corp., No. 20-2152 (D. Minn. Sept. 28, 2022). The use of a generative model to produce the defendant’s work can alternatively support the finding of access through the possibility that the model accessed the original work, given the model’s extensive potential to access vast datasets and the common practice of training models on diverse web-scraped data.
4.2. NAF as a Framework for Copyright Safety
The NAF condition (Vyas et al., 2023) ensures that the generation probability of a generative model is closely aligned with that of a version of the model that has not had access to the copyrighted material. In a canonical form (Vyas et al., 2023; Elkin-Koren et al., 2024), the NAF condition can be understood to require the following:
$P(Z=z|Access=1)\leq e^{\epsilon}P(Z=z|Access=0)$
for any $z$, where $Z$ is a random variable representing the output of the generative model, $Access$ indicates whether the model had access to a copyrighted work, and parameter $\epsilon>0$ determines the strength of theoretical guarantee (the smaller the value of $\epsilon$, the closer the output distribution of the model must be to that of a model trained without access). The output distribution of the AI model must satisfy this inequality with respect to any copyrighted work in the training data. The NAF condition with a specific parameter value $\epsilon$ will be referred to as $\epsilon$-NAF herein. One should note that this formulation is idealized, and in reality, the existence or nonexistence of access may not be clear-cut, as copyright-protected elements can be widely distributed and repeated across training data in ways that make them inseparable from other data.
While the authors suggest that this criterion is related to the concept of substantial similarity in copyright law, they leave open the question of whether this approach could function as a valid defense in court (Vyas et al., 2023). Although the NAF condition has been acknowledged as at least partially effective in reducing the risks of infringement (Elkin-Koren et al., 2024), it has also faced skepticism for not fully aligning with the principles of copyright law (Lee et al., 2024). Here, probabilistic formalism is employed as an aid for precise analysis to reconcile diverging views.
One straightforward way the NAF condition contributes to copyright safety is by addressing the prospective problem of how likely the model is to generate a copyright-infringing output. It can be concluded that this probability is low, given that (i) the model satisfies $\epsilon$-NAF with a small $\epsilon$ and (ii) the model is unlikely to produce outputs highly similar to copyrighted works if it had been trained without access. This trivially follows from standard probabilistic reasoning by viewing the substantial similarity of a generated work to a fixed copyrighted work as a function of the generated work. Specifically, letting $Substantial(Z)$ be a binary function of random variable $Z$ that takes value 1 if the similarity of an output $z$ to the copyrighted work reaches the threshold corresponding to substantial similarity, $\epsilon$-NAF ensures that:
$\begin{split}&P(Substantial(Z)=1|Access=1)\\ &\leq e^{\epsilon}P(Substantial(Z)=1|Access=0).\end{split}$
Due to this inequality, given that $P(Substantial(Z)=1|Access=0)$ is small and $e^{\epsilon}$ is close to 1, corresponding to a small value of $\epsilon$, $P(Substantial(Z)=1|Access=1)$ must also be small. Consequently, the model is unlikely to generate content that is substantially similar to the copyrighted work, and thus unlikely to infringe on its copyright, as its generation would not satisfy the improper appropriation prong.
However, the retrospective problem of determining whether a specific generated output infringes copyright—the issue adjudicators face in court—is more nuanced. Drawing a conclusion based on broad statistical data that generative models satisfying the NAF condition produce substantially similar outputs with low probability, without assessing case-specific details pertaining to the substantial similarity between the specific output and the original work at hand, constitutes a fallacy that must be avoided (Tribe, 1971; Allen and Stein, 2013). The mere rarity of substantially similar outputs generated by such models does not influence the evaluation of substantial similarity between the two specific works, which adjudicators can typically ascertain with certainty by directly examining both works presented in court. Therefore, linking the NAF condition to the improper appropriation prong in this context appears unfounded.
The focus now shifts to examining whether and how the NAF condition affects the determination of the other prong, factual copying, within the context of the retrospective problem. On one hand, there seems to be little justification for expecting the NAF condition to meaningfully influence the finding of copying, given both the presence of access and a fixed level of similarity in the probative sense. Specifically, the theoretical guarantee of the NAF condition do not directly impose constraints on factual copying; instead, it merely dictates a probabilistic relationship between access and generated outputs. Furthermore, as observed in (Lee et al., 2024), the assertion that a similar output could have been produced by copying from another source does not constitute a valid defense under copyright law.
On the other hand, the NAF condition can impact the inference of access based on similarity, as it influences the observable distribution of similarity between generated content and copyrighted works. Random variables as defined in Definition 4.1 is considered to analyze the implications.
Definition 4.1.
$A_{M}\in\{0,1\}$, $S\in\{1,2,\cdots,n\}$, and $EA_{M}\in\{1,2,\cdots,m\}$ are random variables that satisfy:
(1) $P(S=i|A_{M}=1)\leq e^{\epsilon}P(S=i|A_{M}=0),\,\forall i$
(2) $P(A_{M}=1|EA_{M}=k)=\delta_{k},\forall k$
(3) $EA_{M}\perp\!\!\!\perp S|A_{M}$
where $\delta_{k+1}\geq\delta_{k}$.
In the context of AI generation, access can refer to two distinct concepts: access by the defendant (as in traditional cases) and access by the AI model. To clarify that access by the model is at issue, the variable for access, $A_{M}$, and the corresponding evidence, $EA_{M}$, are denoted with the subscript $M$. It is of interest to see how the NAF condition (Definition 4.1 (1)) affects the inference of access given similarity and other relevant evidence, or in other words, how it affects the value of $P(A_{M}=1|S=i,\,EA_{M}=k)$. To study this, some degree of belief, $\delta_{k}$, about the likelihood of access by the AI model given other evidence without the knowledge of the level of similarity is assumed (Definition 4.1 (2)). Analogous to $EA$ in Definition 3.1, $EA_{M}$ directly impacts only the inference of access (Definition 4.1 (3)).
Given this setting, Proposition 4.2 on the effectiveness of the NAF condition can be obtained.
Proposition 4.2.
For any $\epsilon>0$ and $(i,k)$, $P(A_{M}=1|S=i,\,EA_{M}=k)\leq\Gamma(\epsilon,\delta_{k})$ where the random variables are defined as in Definition 4.1 and $\Gamma(\epsilon,\delta):=e^{\epsilon}\delta/(1+\delta(e^{\epsilon}-1))$. $\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$, and satisfies $\delta\leq\Gamma(\epsilon,\delta)$ and $\lim_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\delta$.
The proposition shows that the NAF condition indeed imposes an upper bound on the probability of finding the presence of access to the original work by the model. A lower value of $\epsilon$ enhances the effectiveness of this upper bound, and, in the limit of $\epsilon\rightarrow 0$, the evidentiary value of similarity in inferring access vanishes. In other words, similarity fails to serve as meaningful additional evidence of access, with the value of $P(A_{M}=1|S=i,\,EA_{M}=k)$ converging to $\delta_{k}=P(A_{M}=1|EA_{M}=k)$ for all $i$.
However, the way in which the NAF condition operates in this context is ethically debatable. The capacity to generate content with comparable likelihood—regardless of whether access occurred—impedes the determination of whether a work was produced with access based on the generation outcome. Intuitively, since a generative model satisfying the NAF condition is likely to produce a work with the same level of similarity even without access to the copyrighted work, the inferential value of similarity in establishing access by the AI model is reduced. It is questionable whether such an approach is justified, especially considering the demand for transparency in training data as advocated by researchers (Sag, 2023; Zhang et al., 2024) and reflected in the EU AI Act, [^9] Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 laying down harmonised rules on artificial intelligence and amending Regulations (EC) No 300/2008, (EU) No 167/2013, (EU) No 168/2013, (EU) 2018/858, (EU) 2018/1139 and (EU) 2019/2144 and Directives 2014/90/EU, (EU) 2016/797 and (EU) 2020/1828 (Artificial Intelligence Act), Official Journal of the European Union, L 1689, 12 July 2024, pp. 1–144. which imposes transparency requirements on developers of certain AI systems, mandating that they make available summaries of the training data used. Furthermore, since $\Gamma$ is bounded below by the degree of belief $\delta_{k}$ that the model had access to the specific original work in question, given available evidence of access other than similarity, the upper bound on the degree of belief about access becomes increasingly insignificant as stronger evidence of access to the specific original work by the model is presented (e.g., unequivocal evidence of the work’s inclusion in its training data) or as it becomes common practice to include broader sets of copyrighted materials in the training of AI models in general, both of which can lead to higher values of $\delta_{k}$.
5. Discussion
Probabilistic analysis has been shown to enhance the understanding of copyright jurisprudence and its interaction with emerging technologies. In particular, a mathematical proof demonstrated that the controversial inverse ratio rule holds, provided it is appropriately defined. Furthermore, the NAF condition was examined as a mitigation strategy for the copyright infringement risks posed by generative AI, revealing ethical concerns and limitations in its efficacy when addressing the retrospective problem adjudicators face in court, where they must determine whether infringement has occurred based on available evidence.
There are certain limitations to the analysis presented in this paper. While the analysis is based on a probabilistic view of legal fact-finding, this perspective has faced criticisms (Tribe, 1971; Allen and Stein, 2013), although it continues to receive support (Cheng, 2013; Burtis et al., 2017). The validity of probabilistic approaches in legal fact-finding in general is widely debated in legal literature and falls beyond the scope of this paper, except for the relation between the results concerning the inverse ratio rule and the conjunction problem. Furthermore, the analysis abstracts from certain details involved in AI generation, such as the content of user-provided prompts, which may influence the identification of direct or indirect infringers (Lee et al., 2024) and be important in actual cases.
Several related directions for future work remain to be explored. Although this paper has paid particular attention to the inverse ratio rule—especially compatible with probabilistic reasoning due to its direct connection to the assessment of evidence—the presented approach has broader potential applications within copyright jurisprudence, as many copyright principles can be understood through the lens of evidentiary rules (Lichtman, 2003). Furthermore, while the analysis has addressed the NAF condition, future research should focus on developing alternative copyright risk mitigation strategies beyond the NAF condition and conducting a more comprehensive analysis of how generative AI technology interacts with copyright law.
Acknowledgements.
The author acknowledges financial support from the Simons Foundation Math+X Grant to the University of Pennsylvania.
Appendix
Appendix A The Inverse Ratio Rule with Strict Inequality
Proposition 3.2 was proved by assuming only very general relationships among the $\alpha_{i}$’s as well as among the $\beta_{i,j}$’s. It does not specify, for example, how the value of $\beta_{i,j}$ changes when one index is increased while the other is decreased simultaneously, nor does it address the relative magnitude of the changes in the values of $\alpha_{i}$ and $\beta_{i,j}$. Proposition A.1 demonstrates that, under certain assumptions regarding these details, the inverse ratio rule holds with strict inequality.
The inequality $\alpha_{\mu^{\prime}}/\alpha_{\mu^{\prime}-1}\leq\beta_{\mu^{\prime}-1,j}/\beta_{\mu^{\prime},j^{\prime}}$ assumed in the first part of Proposition A.1 roughly states that the increase in the degree of belief in access resulting from stronger evidence of access (excluding similarity) is large relative to the changes in the degrees of belief in both access and copying that arise from different levels of similarity. On the other hand, the inequality $\beta_{i^{\prime},\kappa^{\prime}}/\beta_{i,\kappa^{\prime}-1}\leq\alpha_{i}/\alpha_{i^{\prime}}$ in the second part of the proposition roughly means that the increase in the degree of belief in copying due to a higher level of similarity is substantial relative to the possible increase (or decrease, in which case the inequality is trivially satisfied) in the degree of belief in access with stronger evidence of access paired with a showing of a lower level of similarity.
Importantly, these relations are local in the sense that they only have to hold for some specific indices for the proposition to apply, and it is not necessary for them to be universally true across all indices $i$ and $j$. It is quite plausible, therefore, that the proposition apply in some actual instances.
Proposition A.1.
Assuming that the random variables are defined as in Definition 3.1, the following strong inverse ratio rule holds for any $\lambda\in(0,1)$:
  • Suppose $j>j^{\prime}$ and $\mu^{\prime}>1$ where $\mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$. Then, $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}<\mu^{\prime}$ if $\alpha_{\mu^{\prime}}/\alpha_{\mu^{\prime}-1}\leq\beta_{\mu^{\prime}-1,j}/\beta_{\mu^{\prime},j^{\prime}}$.
  • Suppose $i>i^{\prime}$ and $\kappa^{\prime}>1$ where $\kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$. Then, $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}<\kappa^{\prime}$ if $\beta_{i^{\prime},\kappa^{\prime}}/\beta_{i,\kappa^{\prime}-1}\leq\alpha_{i}/\alpha_{i^{\prime}}$.
Appendix B Relation to the Conjunction Problem
The conjunction problem is frequently cited to criticize probabilistic approaches to legal fact-finding (Pardo and Allen, 2008). It refers to the issue that when a claim consists of multiple elements that must each be established by evidence, the probability of the conjunction of these elements may not reach a required threshold $\lambda$, even if the probability of each individual element exceeds $\lambda$. More concretely, suppose there are two events, $E_{1}$ and $E_{2}$, that need to be established to make a certain claim. The probability that both occurs is the product of the probabilities of the two elements, $P(E_{1}\cap E_{2})=P(E_{1}|E_{2})P(E_{2})$, which is smaller than the probability of either individual element. Thus, even if both elements have been shown to have probabilities higher than $\lambda$ individually, $P(E_{1}|E_{2}),P(E_{2})>\lambda$, it can be the case that the probability of both requirements being satisfied simultaneously is below the threshold, $P(E_{1}\cap E_{2})<\lambda$. A solution offered by the probabilistic camp is to require that the probability of the conjunction, in addition to the individual probability of each element, exceeds the threshold in order to make an affirmative finding (Nance, 1986; Friedman, 1992). However, critics object to this solution for its counterintuitiveness (Pardo and Allen, 2008).
Such objection may appear to affect the results in this paper regarding the inverse ratio rule, which relies on how the probability of copying being a product of two terms: the probability of copying given the presence of access and the probability of access. However, this is not the case because the situation here differs from those considered in the context of the conjunction problem. Under the Arnstein test, it is not required to prove access to the original work as a separate element from copying. Instead, the plaintiff merely needs to show that the probability of copying reaches the threshold, $P(C=1|S=i,\,EA=j)>\lambda$. As shown in the proof of Proposition 3.1, due to the logical relationship between copying and access (Definition 3.1 (1)), this probability can be expressed as the product of two terms, the probability of copying and that of access, leading to:
$\begin{split}&P(C=1|S=i,\,EA=j)>\lambda\\ &\Leftrightarrow P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)>\lambda.\end{split}$
This product arises merely as a consequence of the logical relationship between copying and access, rather than from adopting the controversial solution that necessitates the probability of the conjunction reaching the threshold, which can be formally described as:
$P(C=1,\,A=1|S=i,\,EA=j)>\lambda.$
Therefore, the argument presented here is not directly influenced by discussions surrounding the conjunction problem.
Appendix C Proofs of Propositions
C.1. Proof of Proposition 3.2
Using the logical relationship between copying and access (Definition 3.1 (1)) as well as $EA$’s independence (Definition 3.1 (4)), for any pair $(i,j)$,
$\begin{split}&P\left(C=1|S=i,\,EA=j\right)\\ &=P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)\\ &\quad+P\left(C=1|A=0,\,S=i\right)P\left(A=0|S=i,\,EA=j\right)\\ &=P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)\\ &\quad+0\cdot P\left(A=0|S=i,\,EA=j\right)\\ &=\alpha_{i}\beta_{i,j}.\end{split}$
The first element of the inverse ratio rule is proved by showing contradiction. Assume $\mu>\mu^{\prime}$ where:
$\begin{gathered}\mu:=\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\\ \mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}.\end{gathered}$
Then, by assumptions about the ordering of the terms of the form $\alpha_{i}$ and $\beta_{i,j}$,
$\begin{split}\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j}\leq\alpha_{\mu}\beta_{\mu,j}\end{split}.$
However, this implies that $\mu^{\prime}$, which is strictly smaller than $\mu$, would have achieved $P(C=1|S=\mu^{\prime},\,EA=j)>\lambda$, contradicting the assumption that $\mu$ was the minimum that achieves this.
The proof for the second element of the inverse ratio rule is similar. Assume $\kappa>\kappa^{\prime}$ where:
$\begin{gathered}\kappa:=\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\\ \kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}.\end{gathered}$
Same as above, the following inequalities hold:
$\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa}$
which results in contradiction with the assumption of the minimality of $\kappa$.
C.2. Proof of Proposition 4.2
By the NAF condition (Definition 4.1 (1)) and conditional independence (Definition 4.1 (3)):
$\begin{split}&P(A_{M}=1|S=i,\,EA_{M}=k)\\ &=\frac{P(S=i|A_{M}=1,\,EA_{M}=k)P(A_{M}=1|EA_{M}=k)}{P(S=i|EA_{M}=k)}\\ &\leq\frac{e^{\epsilon}P(S=i|A_{M}=0,\,EA_{M}=k)P(A_{M}=1|EA_{M}=k)}{P(S=i|EA_{M}=k)}\\ &=\frac{e^{\epsilon}\frac{P(A_{M}=0|S=i,\,EA_{M}=k)P(S=i|EA_{M}=k)}{P(A_{M}=0|EA_{M}=k)}P(A_{M}=1|EA_{M}=k)}{P(S=i|EA_{M}=k)}\\ &=\frac{e^{\epsilon}(1-P(A_{M}=1|S=i,\,EA_{M}=k))P(A_{M}=1|EA_{M}=k)}{1-P(A_{M}=1|EA_{M}=k)}.\end{split}$
By rearranging terms,
$\begin{split}P(A_{M}=1|S=i,\,EA_{M}=k)&\leq\frac{e^{\epsilon}P(A_{M}=1|EA_{M}=k)}{1+P(A_{M}=1|EA_{M}=k)(e^{\epsilon}-1)}\\ &=\frac{e^{\epsilon}\delta_{k}}{1+\delta_{k}(e^{\epsilon}-1)}\\ &=\Gamma(\epsilon,\delta_{k}).\end{split}$
Taking the derivative of $\Gamma(\epsilon,\delta)$ with respect to $\epsilon$ and $\delta$, respectively,
$\begin{split}&\frac{d\Gamma}{d\mathcal{\epsilon}}=\frac{e^{\epsilon}\delta(1-\delta)}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0\\ &\frac{d\Gamma}{d\delta}=\frac{e^{\epsilon}}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0.\end{split}$
Therefore, $\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$.
The inequality $\delta\leq\Gamma(\epsilon,\delta)$ follows from the observation that:
$\Gamma(\epsilon,\delta)-\delta=\frac{\delta(e^{\epsilon}-1)(1-\delta)}{1+\delta(e^{\epsilon}-1)}\geq 0$
where the inequality follows from the fact that each term in the numerator and denominator is non-negative because $1\geq\delta\geq 0$ and $e^{\epsilon}>1$.
Finally, the limit as $\epsilon$ approaches zero can be computed straightforwardly as:
$\lim_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\frac{1\cdot\delta}{1+\delta\cdot 0}=\delta.$
C.3. Proof of Proposition A.1
As in the proof of Proposition 3.2,
$P(C=1|S=i,\,EA=j)=\alpha_{i}\beta_{i,j},\,\forall(i,j).$
For the first part, observe:
$\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}.$
In particular, we have $\lambda<\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}$. This implies $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\mu^{\prime}-1<\mu^{\prime}$, which completes the proof of the first part. Similarly, for the second part, we have:
$\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}-1}.$
This implies $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\kappa^{\prime}-1<\kappa^{\prime}$, which completes the proof.

### v5

@@
 This paper presents a probabilistic approach to analyzing copyright infringement disputes.
-Under this approach, evidentiary principles shaped by case law are formalized in probabilistic terms, allowing for a mathematical examination of issues arising in such disputes.
-The usefulness of this approach is showcased through its application to the “inverse ratio rule”—a controversial legal doctrine adopted by some courts.
+Evidentiary principles shaped by case law are formalized in probabilistic terms, and the “inverse ratio rule”—a controversial legal doctrine adopted by some courts—is examined.
 Although this rule has faced significant criticism, a formal proof demonstrates its validity, provided it is properly defined.
-Furthermore, the paper employs the probabilistic approach to study the copyright safety of generative AI.
+The probabilistic approach is further employed to study the copyright safety of generative AI.
 Specifically, the Near Access-Free (NAF) condition, previously proposed as a strategy for mitigating the heightened copyright infringement risks of generative AI, is evaluated.
-The analysis reveals that, while the NAF condition mitigates some infringement risks, its justifiability and efficacy are questionable in certain contexts.
-These findings illustrate how taking a probabilistic perspective can enhance our understanding of copyright jurisprudence and its interaction with generative AI technology.
+The analysis reveals limitations in its justifiability and efficacy.
 copyright: noneccs: Social and professional topics Copyrightsccs: Computing methodologies Probabilistic reasoning
@@
 This article employs a probabilistic approach to analyze copyright infringement disputes, seeking to bring clarity to its complexities.
-The analysis is carried out by distilling and formalizing the fundamental principles governing copyright disputes as established through case law in the United States, particularly focusing on the requirements a copyright owner must meet to make a prima facie case of copyright infringement in court.
-This body of jurisprudence encompasses not only the substantive elements of copyright law, but also the evidentiary principles that guide how evidence is presented and assessed (Loren and Reese, 2019).
-The theoretical foundation of this work lies in the long-recognized link between probabilistic reasoning and judicial proof (Eggleston, 1989; Cullison, 1969; Broun and Kelly, 1970; Finkelstein and Fairley, 1970; Cohen, 1977; Eggleston, 1978; McCauliff, 1982).
-Specifically, the legal fact-finding process can be understood to involve assessment of the probabilities of allegations being true or false based on the available evidence (Stein, 2005a; Cheng, 2013).
-Building on this link, core evidentiary principles in copyright jurisprudence can be formalized in probabilistic terms, allowing for a mathematical analysis of a controversial legal doctrine, the “inverse ratio rule.” This rule governs the relationship between two critical pieces of evidence in copyright infringement cases—access and similarity to the original work.
-Although the validity of this rule has been the subject of considerable debate (Aronoff, 2007; Balganesh and Menell, 2022), a formal proof is established here, demonstrating that it holds under natural assumptions when its content and scope are carefully delineated.
-Furthermore, this approach facilitates the analysis of copyright infringement risks introduced by technological advancements and the identification of effective mitigation strategies.
-Specifically, the probabilistic approach is utilized to evaluate the efficacy of the Near Access-Free (NAF) condition (Vyas et al., 2023), which has been proposed to mitigate the risk of copyright infringement posed by generative AI.
-This paper identifies extensive access to copyrighted materials by generative models as a significant source of heightened infringement risks.
-While the NAF condition addresses this issue to some extent, the analysis reveals that it does so partly on ethically questionable grounds—namely, by reducing transparency regarding the materials to which generative AI had access during training—and that its effectiveness in mitigating infringement risks remains limited under certain situations.
-2. Related Work
-Existing research has investigated the relationship between copyright law and probabilistic thinking, specifically concerning how the probability of creation relates to concepts of creativity and originality within the context of copyright law (Byron, 2006; Khong, 2007; Vermont, 2012; Hacohen and Elkin-Koren, 2024; Chiba-Okabe and Su, 2024).
-However, there has been little discussion on the application of probabilistic approaches in analyzing the evidentiary rules of copyright infringement disputes.
-While a previous study has alluded to the connection between the probabilistic nature of legal fact-finding and the inverse ratio rule (Balganesh and Menell, 2022), no formal probabilistic analysis of the rule has been undertaken, to the best of the author’s knowledge.
-This gap is addressed by employing a probabilistic approach to formalize the evidentiary principles governing copyright disputes and mathematically analyze the inverse ratio rule.
-In response to rapid advancements in generative AI technology, a substantial body of literature has emerged, addressing potential copyright risks and defenses available to developers (Sag, 2023; Henderson et al., 2023; Lee et al., 2024).
-These works discuss various legal challenges, including the potential for copyright infringement arising from AI-generated content, with some studies analyzing the Near Access-Free (NAF) condition as a mitigation strategy (Lee et al., 2024; Elkin-Koren et al., 2024).
-This paper contributes to the discourse by providing a probabilistic analysis of legal issues surrounding the NAF condition, particularly through the lens of inference based on evidence in copyright infringement disputes.
-3. Analysis of Copyright Disputes
-3.1. Essential Evidentiary Principles in Copyright Disputes
+Based on the view that the legal fact-finding process can be understood to involve the evaluation of the probabilities that the allegations are true or false based on available evidence (Cullison, 1969; Finkelstein and Fairley, 1970; Stein, 2005), the core evidentiary principles of copyright jurisprudence are formalized in probabilistic terms.
+This allows for a mathematical analysis of a controversial copyright legal doctrine, the “inverse ratio rule,” whose validity has been the subject of considerable debate (Aronoff, 2007; Balganesh and Menell, 2022).
+While a previous study (Balganesh and Menell, 2022) notes a link between the probabilistic nature of legal fact‑finding and the inverse ratio rule, no formal probabilistic treatment of this rule appears in the literature.
+This paper fills that gap by offering a formal analysis of the inverse ratio rule, demonstrating that the rule must hold under natural assumptions when its content and scope are carefully delineated.
+The same approach is further utilized to evaluate the Near Access-Free (NAF) condition (Vyas et al., 2023), which has been proposed to mitigate the copyright infringement risks posed by generative AI.
+While the authors who introduced the NAF condition left its precise legal characterization an open question (Vyas et al., 2023), subsequent studies have examined the NAF condition from a legal perspective (Lee et al., 2024; Elkin-Koren et al., 2024).
+This paper advances the discussion by revealing that, in certain contexts, the NAF condition operates in an ethically questionable way—reducing transparency about the materials generative AI accessed during training—and that its effectiveness remains limited.
+2. Probabilistic Analysis of Copyright Disputes
+2.1. Formalization of Evidentiary Principles
 To prevail in a copyright infringement dispute, the copyright owner, as the plaintiff, must demonstrate two elements under the traditional Arnstein test: “(a) that defendant copied from plaintiff’s copyrighted work and (b) that the copying (assuming it to be proved) went so far as to constitute improper appropriation.” [^1] Arnstein v.
@@
 These two requirements are often referred to simply as “factual copying” and “improper appropriation,” respectively (Balganesh and Menell, 2022).
-Although several defenses, including fair use, may be available to the defendant in some cases even if these elements are established (Loren and Reese, 2019), these defenses are not considered in this paper to maintain focus on the essential structure of copyright infringement cases.
 Direct evidence of factual copying includes admission by the defendant or an eyewitness account of the replication or appropriative behavior, but such evidence is hard to obtain.
@@
 Courts generally hold that proof of access combined with a certain level of similarity, that is probative of copying, is sufficient to establish factual copying (Latman, 1990; Loren and Reese, 2019).
-Although both access and similarity support the inference of factual copying, simply having proof of access does not suffice to establish factual copying when there is no similarity between the two works.
-[^3] Arnstein, 154 F.2d at 468.
-Proof of access pertains to the antecedent of factual copying, in the sense that the defendant having had access to the original work is a necessary condition for engaging in copying (Balganesh and Menell, 2022; Latman, 1990).
-In contrast, similarity emerges as a result of copying (Balganesh and Menell, 2022; Latman, 1990), and when the similarity between the two works is so pronounced as to be deemed “striking,” this alone suffices to establish that copying has occurred.
-[^4] Id.
-at 468-69.
+Access pertains to the antecedent of factual copying, in the sense that the defendant having had access to the original work is a necessary condition for engaging in copying (Balganesh and Menell, 2022; Latman, 1990).
+In contrast, similarity emerges as a result of copying (Balganesh and Menell, 2022; Latman, 1990).
+When the similarity between the two works is so pronounced as to be deemed “striking,” this alone suffices to establish that copying has occurred.
+[^3] Arnstein, 154 F.2d at 468-69.
 In such cases, it can be inferred that the defendant had access to the original work based on the high degree of similarity.
@@
 The second requirement, improper appropriation, is established by demonstrating that the extent of copying reaches the threshold at which the similarity between the original work and the defendant’s work can be deemed “substantial” in degree.
-Determination of this requirement involves a normative judgment, taking into account both quantitative and qualitative aspects of copying (Gorman, 2006; Balganesh, 2012).
+The determination of this requirement involves a normative judgment, taking into account both quantitative and qualitative aspects of copying (Gorman, 2006; Balganesh, 2012).
 Thus, similarity serves a dual role: It supports the finding of factual copying and acts as an independent criterion for establishing improper appropriation (Latman, 1990; Balganesh and Menell, 2022), though the sense of similarity differs in each context, even if there may be some overlap (Loren and Reese, 2019).
-3.2. Formalization of Evidentiary Principles
-Legal fact-finding is a reasoning process that involves assessing and weighing available evidence to reach a decision amid uncertainty.
-This process may be appropriately characterized by probabilistic reasoning, wherein adjudicators employ common sense and logic, drawing on knowledge of the frequencies of occurrences of certain events and patterns derived from experience to evaluate the likelihood of the particular circumstances in question (Stein, 2005b).
-Although the absolute truth typically cannot be attained, the final decision in most civil cases is made based on the standard of “preponderance of the evidence.” This standard requires the party bearing the burden of proof to demonstrate that their claim is more likely true than not.
-From a probabilistic perspective, it is often understood as requiring a demonstration of greater than 50% likelihood that the claim is true (McCauliff, 1982; Stein, 2005a), particularly in binary situations where the question is whether an assertion is true or false (Cheng, 2013).
-In criminal cases, the relevant standard is “proof beyond a reasonable doubt,” which is commonly understood to require a significantly higher probability (McCauliff, 1982).
-Probabilistic approaches to fact-finding have been extensively explored in the literature over many years (Eggleston, 1989; Cullison, 1969; Broun and Kelly, 1970; Finkelstein and Fairley, 1970; Cohen, 1977; Eggleston, 1978; McCauliff, 1982) and, despite considerable criticisms (Tribe, 1971; Allen and Stein, 2013), have remained influential, even if primarily at a conceptual level.
-Building upon this foundation, the essential structure of copyright disputes can be formalized by representing adjudicators’ degree of belief in terms of subjective probabilities, utilizing a discrete probability space $(\Omega,\mathcal{F},P)$ and random variables defined on this space.
+In court proceedings, adjudicators draw on knowledge of the frequencies of occurrences of certain events and patterns derived from experience to evaluate the likelihood of the particular circumstances in question (Stein, 2005).
+Although the absolute truth typically cannot be attained, the final decision in most civil cases is made based on the “preponderance of evidence.” This standard requires the party bearing the burden of proof to demonstrate that their claim is more likely true than not.
+From a probabilistic perspective, it is often understood as requiring a demonstration of a likelihood greater than 50% that the claim is true (McCauliff, 1982; Stein, 2005), particularly in binary situations where the question is whether an assertion is true or false (Cheng, 2013).
+Building on this view and considering a discrete probability space $(\Omega,\mathcal{F},P)$, essential copyright evidentiary principles can be formalized in probabilistic terms.
 The condition for copyright infringement to exist can be expressed using random variables as $Copy=1\wedge Substantial=1$.
-Here, $Copy$ is a binary random variable representing the presence ($Copy=1$) or non-presence ($Copy=0$) of factual copying of the plaintiff’s copyrighted work during the creation of the defendant’s work.
-Similarly, $Substantial$ represents the presence or non-presence of substantial similarity between the two works.
+Here, $Copy$ is a binary random variable representing the presence ($Copy=1$) or absence ($Copy=0$) of factual copying of the plaintiff’s copyrighted work during the creation of the defendant’s work.
+Similarly, $Substantial$ represents the presence or absence of substantial similarity between the two works.
 Typically, adjudicators can determine whether substantial similarity exists, or the truth value of $Substantial$, with certainty by directly comparing the original work to the defendant’s work presented at trial.
-In contrast, the absolute truth about whether the defendant engaged in factual copying remains unknown during the trial, necessitating a probabilistic approach to account for this uncertainty.
-The plaintiff must present sufficient evidence for the adjudicators to conclude that the probability of $Copy=1$ exceeds a certain threshold, $\lambda$, which could be 0.5, aligning with the preponderance of evidence standard.
-The case law establishing that probative similarity combined with access typically suffices to prove copying can be represented in probabilistic terms as follows:
-$\begin{gathered}P(Copy=1|Access=1,\,Probative=1)=\eta\\ \textrm{such that}\,\,\eta>\lambda\end{gathered}$
-where $Access$ and $Probative$ are binary random variables representing whether access by the defendant to the plaintiff’s original work was present and whether the original and defendant’s works are probatively similar, respectively.
+In contrast, the absolute truth about whether the defendant engaged in factual copying remains unknown during the trial, and the plaintiff must present sufficient evidence for the adjudicators to conclude that the probability of $Copy=1$ exceeds a certain threshold, $\lambda$, corresponding to the preponderance of evidence standard.
+The case law establishing that probative similarity combined with access typically suffices to prove copying can be represented in probabilistic terms as $P(Copy=1|Access=1,\,Probative=1)=\eta$ where $\eta>\lambda$.
+$Access$ and $Probative$ are binary random variables representing whether access by the defendant to the plaintiff’s original work was present and whether the original and defendant’s works are probatively similar, respectively.
 Suppose the adjudicators learn from the evidence presented during the trial that the defendant indeed had access to the original work ($Access=1$) and that the defendant’s work is probatively similar to the original work ($Probative=1$).
-This results in the adjudicators’ degree of belief that copying occurred ($Copy=1$) surpassing the threshold $\lambda$ (since $\eta>\lambda$), thereby the standard of proof being met, as desired.
-This approach can accommodate the logical relationship between factual copying and access, specifically that the defendant must have had access to the original work in order to copy from it.
-This relationship can be represented as:
-$P(Copy=1|Access=0)=0.$
-As a result of imposing this restriction on $Copy$, if it is known with certainty that the defendant did not have access to the original work, the degree of belief that the defendant copied the original work must be zero, as expected.
-A further step can be taken to represent the evidentiary rule that striking similarity is typically sufficient to establish copying as follows:
-$\begin{gathered}P(Copy=1|Access=1,\,Striking=1)=\rho\\ P(Access=1|Striking=1)=\sigma\\ \textrm{such that}\,\,\rho\sigma>\lambda\end{gathered}$
-where $Striking$ is a binary random variable corresponding to the presence or non-presence of striking similarity.
-Assuming that the adjudicators learn during the trial that $Striking=1$ with certainty, taking into account the logical relationship between access and copying, the degree of belief about copying can be computed as:
-$\begin{split}&P\left(Copy=1|Striking=1\right)\\ &=P\left(Copy=1|Striking=1,\,Access=1\right)P\left(Access=1|Striking=1\right)\\ &\,\,+P\left(Copy=1|Striking=1,\,Access=0\right)P\left(Access=0|Striking=1\right)\\ &=P\left(Copy=1|Striking=1,\,Access=1\right)P\left(Access=1|Striking=1\right)\\ &\,\,+0\cdot P\left(Access=0|Striking=1\right)\\ &=\rho\sigma>\lambda.\end{split}$
-Therefore, the degree of belief of the adjudicators reaches the threshold corresponding to the applicable standard of proof, allowing them to conclude that copying has occurred.
+This results in the adjudicators’ degree of belief that copying occurred ($Copy=1$) surpassing the threshold $\lambda$ (since $\eta>\lambda$), thus meeting the standard of proof.
+This approach can accommodate the aforementioned logical relationship between factual copying and access, which can be represented as $P(Copy=1|Access=0)=0$.
+A further step can be taken to represent the evidentiary rule that striking similarity is typically sufficient to establish copying by stipulating $P(Copy=1|Access=1,Striking=1)=\rho$ and $P(Access=1|Striking=1)=\sigma$ where $\rho\sigma>\lambda$.
+$Striking$ is a binary random variable corresponding to the presence or absence of striking similarity.
+Assuming that the adjudicators learn during the trial that $Striking=1$ with certainty, taking into account the logical relationship between access and copying, the degree of belief about copying exceeds the threshold, as desired: $P\left(Copy=1|Striking=1\right)=P\left(Copy=1|Striking=1,Access=1\right)P\left(Access=1|Striking=1\right)+0\cdot P\left(Access=0|Striking=1\right)=\rho\sigma>\lambda$.
 2.2. The Inverse Ratio Rule
-The so-called “inverse ratio rule,” adopted by some courts, dictates a specific relationship between proof of access and similarity: stronger proof of access lowers the level of similarity required to establish copyright infringement, while a higher level of similarity reduces the burden of proof regarding access (Aronoff, 2007; Balganesh and Menell, 2022).
+According to the inverse ratio rule, a stronger proof of access lowers the level of similarity required to establish copyright infringement, while a higher level of similarity reduces the burden of proof regarding access (Aronoff, 2007; Balganesh and Menell, 2022).
 Although the rule has been applied for nearly a century, serious objections have been raised to its application (Aronoff, 2007).
-The controversial status of the inverse ratio rule gained renewed attention with its recent abrogation by the Ninth Circuit Court, [^6] Skidmore v.
+The controversial status of the inverse ratio rule gained renewed attention with its recent abrogation by the Ninth Circuit Court.
+[^5] Skidmore v.
 Led Zeppelin, 952 F.3d 1051, 1066-69 (9th Cir.
 2020) (en banc).
-while it remains in use in some other courts (Balganesh and Menell, 2022).
-The controversy surrounding the inverse ratio rule partly arises from a likely inadvertent divergence in the understanding of its scope and content.
-Some courts have applied the inverse ratio rule in a way that strong evidence of access lowers the level of similarity required to establish the improper appropriation prong necessary for proving infringement (Balganesh and Menell, 2022; Aronoff, 2007).
-However, there appears to be little justification for the inverse ratio rule operating in this manner (Balganesh and Menell, 2022).
-Another, more promising formulation of the inverse ratio rule pertains to the factual copying prong, where the rule adjusts the degree of similarity or the strength of evidence of access required to establish factual copying.
-Specifically, when strong evidence of access is present, a lower degree of similarity is sufficient to prove factual copying, whereas weak evidence of access necessitates a higher level of similarity.
-Conversely, a high degree of similarity lessens the evidence of access required, and vice versa.
-Proponents argue that the inverse ratio rule of this form follows from general principles of evidential inference (Balganesh and Menell, 2022).
-However, critics contend that even the inverse ration rule formulated in this way is flawed, as access is merely a necessary condition for copying, and better evidence of access does not increase the probability of copying (Aronoff, 2007).
-This formulation of the inverse ratio rule was also the one rejected by the Ninth Circuit Court.
-[^7] Id.
-at 1066-68.
-The analysis here precisely describes the content of the inverse ratio rule as it pertains to the factual copying element in formal language and proves that the rule, defined in this way, must hold.
-The proof relies solely on assumptions that are naturally deduced from well-established jurisprudence.
-In particular, we consider a set of random variables that encapsulates the evidentiary principles discussed in the preceding section in a more generalized form (Definition 3.1).
+Some courts have applied the inverse ratio rule to ease the substantial similarity threshold for proving improper appropriation when access is strongly established (Balganesh and Menell, 2022; Aronoff, 2007), but such application of the rule lacks clear justification (Balganesh and Menell, 2022).
+A more promising formulation of the inverse ratio rule concerns the factual copying prong: Strong evidence of access lowers the similarity needed to prove factual copying, while weak evidence of access requires greater similarity; conversely, a high degree of similarity can compensate for limited evidence of access (Balganesh and Menell, 2022).
+In fact, this version of the inverse ratio rule can be proved by considering a set of random variables (Definition 2.1) that encapsulates the evidentiary principles discussed in the preceding section in a more generalized form.
 Definition 2.1.
@@
 $C$ and $A$ are abbreviations of $Copy$ and $Access$, respectively, from earlier discussions, which enjoy a logical relationship (Definition 2.1 (1)).
-Possible values of $S$ represent varying degrees of similarity between the original and defendant’s works in the probative sense.
-As reflected in the well-established evidentiary rule that copying can be established by demonstrating both access and probative similarity, similarity serves as significant evidence in proving factual copying (Definition 3.1 (2)).
-It is reasonable, at least in principle (assuming no other specific circumstances suggest otherwise), to think that there is a monotonic relationship among the inferential values of different degrees of similarity; that is, a higher degree of similarity more strongly infers factual copying, so that $\alpha_{i+1}\geq\alpha_{i}$ holds.
+The value of $S$ represents the level of similarity in the probative sense, which serves as evidence to prove factual copying (Definition 2.1 (2)).
+It is reasonable, at least in principle (assuming that no other specific circumstances suggest otherwise), to think that there is a monotonic relationship among the inferential values of different degrees of similarity; that is, a higher degree of similarity more strongly infers factual copying, so $\alpha_{i+1}\geq\alpha_{i}$ holds.
 Similarity also has evidentiary value in establishing access to the original work (Definition 2.1 (4)), as demonstrated by the evidentiary rule regarding striking similarity.
-Even critics of the inverse ratio rule acknowledge the evidentiary value of similarity for establishing access, though they suggest that the strength of the inference may be weaker than assumed by others (Aronoff, 2007).
 It is natural to posit that such evidentiary value for access increases monotonically with similarity in principle, leading to the assumption that $\beta_{i+1,j}\geq\beta_{i,j}$ for all $i$.
-Variable $EA$ is introduced to capture the interaction between levels of similarity and evidence of access (other than similarity itself).
+Variable $EA$ is introduced to capture the interaction between similarity and evidence of access (other than similarity itself).
 The values it can take correspond to varying degrees of the strength of evidence of access presented in court (Definition 2.1 (3)).
@@
 Conversely, the fact that the original work was kept private, with little chance of the defendant encountering it, would imply a low likelihood of access.
-Such evidence is assumed to have direct impact only on the inference of access and is independent of other variables given $A$ (Definition 3.1 (4)).
-The differing strengths of evidence naturally create an ordered variation in the inferential value of access, captured by the relationship $\beta_{i,j+1}\geq\beta_{i,j}$ for all $j$.
+Such evidence is assumed to directly impact only the inference of access as specified in Definition 2.1 (4), which states that $C$ and $S$ are independent of $EA$ conditional on $A$.
+The differing strengths of evidence naturally create an ordered variation in the inferential value, captured by the relationship $\beta_{i,j+1}\geq\beta_{i,j}$ for all $j$.
 During court proceedings, adjudicators learn both the level of similarity between the works and the strength of evidence of access presented, corresponding to a pair of values $(i,j)$.
-Then, the degree of belief of the adjudicators about the presence of factual copying, $C=1$, are formed given the knowledge of $S=i$ and $EA=j$.
-Under these assumptions, the inverse ratio rule as it pertains to factual copying can now be precisely articulated in formal language and its correctness proved, as summarized in Proposition 3.2 (see Appendix C for proofs).
+Then, the degree of belief of the adjudicators about the presence of factual copying, $C=1$, is formed given the knowledge of $S=i$ and $EA=j$.
+Under these assumptions, the inverse ratio rule as it pertains to factual copying can now be precisely articulated in formal language and its correctness proved, as summarized in Proposition 2.2 (see Appendix A for a stronger version of the inverse ratio rule that holds with strict inequalities).
+Proofs are shown in the Appendix.
 Proposition 2.2.
-The following inverse ratio rule holds for any $\lambda\in(0,1)$:
-  • If $j\geq j^{\prime}$, $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$
-  • If $i\geq i^{\prime}$, $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$
-where the random variables are defined as in Definition 3.1.
-The fact that access constitutes a necessary condition for copying does not imply that the strength of the evidence of access cannot affect the inference of copying, as critics of the inverse ratio rule argue.
-Stronger evidence of access can increase the likelihood of copying because the likelihood of copying is the product of two factors: (i) the likelihood of copying given the presence of access and the level of similarity, and (ii) the likelihood of access given the level of similarity and other evidence of access.
-Thus, stronger evidence of access leads to an increased likelihood of copying by elevating the value of the second factor in the product.
-Although Proposition 3.2 shows the inverse ratio rule only with weak inequalities, which can be proved with minimal assumptions, the interplay between these two factors allows the inverse ratio rule to hold in an even stronger sense, with strict inequalities, in certain specific yet highly plausible circumstances (see Appendix A for further details).
-Furthermore, although the conclusion that the inverse ratio rule holds true is accepted, the analysis indicates that the reasoning posited by its proponents does not fully capture the intricacies involved.
-The argument previously put forward to support the inverse ratio rule contends that two pieces of relatively weak evidence—pertaining to access and similarity, respectively—possess strong inferential value when combined (Balganesh and Menell, 2022).
-However, the reasoning of the results concerning the inverse ratio rule demonstrates that the level of similarity and the strength of evidence on access interact also in a complementary manner, rather than merely cumulatively, due to their multiplicative effect on the probability of copying.
-Specifically, a higher degree of similarity increases factor (i), and even when combined with a relatively lower factor (ii) resulting from weaker evidence of access, the product of these probabilities can still surpass the necessary threshold for establishing factual copying, and vice versa.
-Since the product of two probabilities is always smaller than either individual factor, combining small factors—stemming from weak evidence in both similarity and access—cannot yield a stronger inference on its own.
-This demonstrates that the cumulative inference logic is insufficient to fully explain the interaction between evidence of access and similarity.
-One might notice the product of the two probability terms is reminiscent of the so-called conjunction problem that has been a point of debate between legal scholars who support probabilistic views of legal fact-finding and those who are skeptical about these views.
-However, as elaborated in Appendix B, the situation at hand is different from typical scenarios that are considered in the context of the conjunction problem, and discussions surrounding the problem does not affect the analysis of the inverse ratio rule.
-4. Copyright Safety for Generative AI
-4.1. Evidentiary Considerations in Copyright Disputes Involving AI-Generated Works
+The following inverse ratio rule holds for any $\lambda\in(0,1)$: If $j\geq j^{\prime}$, then $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$.
+Similarly, if $i\geq i^{\prime}$, then $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$.
+Even though access is merely a necessary condition for copying, the strength of its evidence still affects the inference of copying by lowering the required similarity, despite critics’ argument that the inverse ratio rule improperly lowers the similarity threshold (Aronoff, 2007).
+This is because the likelihood of copying is the product of two factors, (i) the likelihood of copying given the presence of access and the level of similarity and (ii) the likelihood of access given the level of similarity and other evidence of access, and stronger (weaker) evidence of access enlarges (dwindles) the second factor.
+Furthermore, the justification offered by the proponents of the inverse ratio rule is clarified in this paper; it is argued that moderate evidence of access and similarity can jointly support an inference of copying (Balganesh and Menell, 2022).
+However, since the product of two probabilities is no greater than either individual factor, combining small factors, stemming from weak evidence in both similarity and access, cannot yield a stronger inference.
+Instead, similarity and access work complementarily rather than cumulatively: Stronger evidence of access raises factor (ii), and even when factor (i) is modest due to a somewhat low level of similarity, their product can still clear the threshold for factual copying, and vice versa.
+3. Near Access-Freeness as a Framework for Copyright Safety of Generative AI
 When generative AI is used to create content, it is typically offered as a service by the AI developer and used by a user.
@@
 In most cases, access and similarity will likely serve as the primary evidence of infringement, much like in traditional copyright disputes (Lee et al., 2024).
-While the creative processes of generative models differ markedly from human creativity, it appears that similar evidentiary principles are likely to apply even when generative AI is involved in the creative processes.
-At a minimum, it is generally recognized that copying, even for generative AI outputs, requires some form of access to the original work, which may be demonstrated by its inclusion in the AI model’s training data (Zirpoli, 2023; Lee et al., 2024; Elkin-Koren et al., 2024).
-Moreover, studies have shown that exposure to certain types of content during training leads to a higher likelihood of these models reproducing or closely imitating that content (Carlini et al., 2021, 2023; Hacohen et al., 2024), suggesting that inclusion of specific data correlates with a greater probability of producing similar outputs through copying.
-This echoes the evidentiary framework applicable to traditional copyright cases, where greater similarity between works strengthens the inference of access and copying.
-A distinctive feature of generative models, setting them apart from human artists and traditional creative tools, is their capacity to be trained on vast and diverse datasets.
-Training datasets are often compiled by scraping content from the Internet, including copyrighted material (Lee et al., 2024).
-Furthermore, direct inclusion of copyrighted material in the training data is not even necessary for latent diffusion models to produce images similar to copyrighted works, as these models can reproduce content from other images that retain similar latent information (Lu et al., 2024), which exacerbates the problem.
-These characteristics suggest a high likelihood that generative models have what can be construed as “access” to copyrighted works.
-As a result, knowing that generative AI was used to create the defendant’s work would likely lead to a stronger belief in the presence of access.
-For example, consider a scenario where the original work was posted online, and the defendant’s work was created by an AI model trained on data scraped from the Internet.
-Even in the Internet age, simply posting a work online does not automatically establish access by the defendant.
-Courts require more than just online availability; factors like the work’s visibility or significant dissemination are needed to strengthen the inference of access.
-[^8] See generally Cooley v.
-Target Corp., No.
-20-2152 (D.
-Minn.
-Sept.
-28, 2022).
-The use of a generative model to produce the defendant’s work can alternatively support the finding of access through the possibility that the model accessed the original work, given the model’s extensive potential to access vast datasets and the common practice of training models on diverse web-scraped data.
-4.2. NAF as a Framework for Copyright Safety
-The NAF condition (Vyas et al., 2023) ensures that the generation probability of a generative model is closely aligned with that of a version of the model that has not had access to the copyrighted material.
-In a canonical form (Vyas et al., 2023; Elkin-Koren et al., 2024), the NAF condition can be understood to require the following:
-$P(Z=z|Access=1)\leq e^{\epsilon}P(Z=z|Access=0)$
-for any $z$, where $Z$ is a random variable representing the output of the generative model, $Access$ indicates whether the model had access to a copyrighted work, and parameter $\epsilon>0$ determines the strength of theoretical guarantee (the smaller the value of $\epsilon$, the closer the output distribution of the model must be to that of a model trained without access).
+The inclusion of a work in a model’s training corpus is often viewed as sufficient evidence of the access element required to establish copying (Lee et al., 2024; Elkin-Koren et al., 2024).
+Near Access-Freeness (Vyas et al., 2023), which aims to mitigate copyright infringement risks, ensures that the generation probability of a generative model is closely aligned with that of a version of the model that has not had access to copyrighted data.
+In a canonical form (Vyas et al., 2023; Elkin-Koren et al., 2024), the NAF condition can be understood to impose the inequality $P(Z=z|Access=1)\leq e^{\epsilon}P(Z=z|Access=0)$ for any $z$, where $Z$ is a random variable representing the output of the generative model, $Access$ indicates whether the model had access to a copyrighted work, and parameter $\epsilon>0$ determines the strength of theoretical guarantee.
 The output distribution of the AI model must satisfy this inequality with respect to any copyrighted work in the training data.
+In words, if a model meets the NAF condition with some $\epsilon$, having access to the original work during training can raise the probability of generating any particular output by at most a factor of $e^{\epsilon}$.
+For instance, a model trained on Snoopy images is at most $e^{\epsilon}$ times more likely to produce any image—including one that resembles Snoopy—than a version of the model that never had access to Snoopy images in training.
 The NAF condition with a specific parameter value $\epsilon$ will be referred to as $\epsilon$-NAF herein.
-One should note that this formulation is idealized, and in reality, the existence or nonexistence of access may not be clear-cut, as copyright-protected elements can be widely distributed and repeated across training data in ways that make them inseparable from other data.
-While the authors suggest that this criterion is related to the concept of substantial similarity in copyright law, they leave open the question of whether this approach could function as a valid defense in court (Vyas et al., 2023).
-Although the NAF condition has been acknowledged as at least partially effective in reducing the risks of infringement (Elkin-Koren et al., 2024), it has also faced skepticism for not fully aligning with the principles of copyright law (Lee et al., 2024).
-Here, probabilistic formalism is employed as an aid for precise analysis to reconcile diverging views.
+The authors who proposed the NAF condition leave open the question of whether this approach could function as a valid defense in court (Vyas et al., 2023).
+Although the NAF condition has been acknowledged to be at least partially effective in reducing the risks of infringement (Elkin-Koren et al., 2024), it has also faced skepticism for not fully aligning with the principles of copyright law (Lee et al., 2024).
+Here, a probabilistic formalism is employed as an aid for precise analysis to reconcile divergent views.
 One straightforward way the NAF condition contributes to copyright safety is by addressing the prospective problem of how likely the model is to generate a copyright-infringing output.
@@
 This trivially follows from standard probabilistic reasoning by viewing the substantial similarity of a generated work to a fixed copyrighted work as a function of the generated work.
-Specifically, letting $Substantial(Z)$ be a binary function of random variable $Z$ that takes value 1 if the similarity of an output $z$ to the copyrighted work reaches the threshold corresponding to substantial similarity, $\epsilon$-NAF ensures that:
-$\begin{split}&P(Substantial(Z)=1|Access=1)\\ &\leq e^{\epsilon}P(Substantial(Z)=1|Access=0).\end{split}$
+Specifically, letting $Substantial(Z)$ be a binary function of random variable $Z$ that takes value 1 if the similarity of an output $z$ to the copyrighted work reaches the threshold corresponding to substantial similarity, $\epsilon$-NAF ensures that: $P(Substantial(Z)=1|Access=1)\leq e^{\epsilon}P(Substantial(Z)=1|Access=0)$.
 Due to this inequality, given that $P(Substantial(Z)=1|Access=0)$ is small and $e^{\epsilon}$ is close to 1, corresponding to a small value of $\epsilon$, $P(Substantial(Z)=1|Access=1)$ must also be small.
-Consequently, the model is unlikely to generate content that is substantially similar to the copyrighted work, and thus unlikely to infringe on its copyright, as its generation would not satisfy the improper appropriation prong.
+Consequently, the model is unlikely to produce substantially similar content or infringe copyright, as it would not meet the improper appropriation prong.
 However, the retrospective problem of determining whether a specific generated output infringes copyright—the issue adjudicators face in court—is more nuanced.
@@
 On one hand, there seems to be little justification for expecting the NAF condition to meaningfully influence the finding of copying, given both the presence of access and a fixed level of similarity in the probative sense.
-Specifically, the theoretical guarantee of the NAF condition do not directly impose constraints on factual copying; instead, it merely dictates a probabilistic relationship between access and generated outputs.
+Specifically, the theoretical guarantee of the NAF condition does not directly impose constraints on factual copying; instead, it merely dictates a probabilistic relationship between access and generated outputs.
 Furthermore, as observed in (Lee et al., 2024), the assertion that a similar output could have been produced by copying from another source does not constitute a valid defense under copyright law.
 On the other hand, the NAF condition can impact the inference of access based on similarity, as it influences the observable distribution of similarity between generated content and copyrighted works.
-Random variables as defined in Definition 4.1 is considered to analyze the implications.
+This becomes evident through a formal analysis employing the random variables as defined in Definition 3.1.
 Definition 3.1.
-$A_{M}\in\{0,1\}$, $S\in\{1,2,\cdots,n\}$, and $EA_{M}\in\{1,2,\cdots,m\}$ are random variables that satisfy:
+$A_{M}\in\{0,1\}$ and $EA_{M}\in\{1,2,\cdots,m\}$ are random variables that satisfy:
 (1) $P(S=i|A_{M}=1)\leq e^{\epsilon}P(S=i|A_{M}=0),\,\forall i$
@@
 where $\delta_{k+1}\geq\delta_{k}$.
-In the context of AI generation, access can refer to two distinct concepts: access by the defendant (as in traditional cases) and access by the AI model.
+In the context of AI generation, access can refer to access by the user or access by the AI model.
 To clarify that access by the model is at issue, the variable for access, $A_{M}$, and the corresponding evidence, $EA_{M}$, are denoted with the subscript $M$.
-It is of interest to see how the NAF condition (Definition 4.1 (1)) affects the inference of access given similarity and other relevant evidence, or in other words, how it affects the value of $P(A_{M}=1|S=i,\,EA_{M}=k)$.
-To study this, some degree of belief, $\delta_{k}$, about the likelihood of access by the AI model given other evidence without the knowledge of the level of similarity is assumed (Definition 4.1 (2)).
+The NAF condition (Definition 3.1 (1)) and some degree of belief, $\delta_{k}$, about the likelihood of access by the AI model given available evidence (other than similarity) of access (Definition 3.1 (2)) are assumed.
 Analogous to $EA$ in Definition 2.1, $EA_{M}$ directly impacts only the inference of access (Definition 3.1 (3)).
-Given this setting, Proposition 4.2 on the effectiveness of the NAF condition can be obtained.
 Proposition 3.2.
-For any $\epsilon>0$ and $(i,k)$, $P(A_{M}=1|S=i,\,EA_{M}=k)\leq\Gamma(\epsilon,\delta_{k})$ where the random variables are defined as in Definition 4.1 and $\Gamma(\epsilon,\delta):=e^{\epsilon}\delta/(1+\delta(e^{\epsilon}-1))$.
-$\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$, and satisfies $\delta\leq\Gamma(\epsilon,\delta)$ and $\lim_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\delta$.
-The proposition shows that the NAF condition indeed imposes an upper bound on the probability of finding the presence of access to the original work by the model.
+For any $\epsilon>0$ and $(i,k)$, $P(A_{M}=1|S=i,\,EA_{M}=k)\leq\Gamma(\epsilon,\delta_{k})$ where $\Gamma(\epsilon,\delta):=\frac{e^{\epsilon}\delta}{1+\delta(e^{\epsilon}-1)}$.
+$\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$, and satisfies $\delta\leq\Gamma(\epsilon,\delta)$ and $\lim\limits_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\delta$.
+Proposition 3.2 shows that the NAF condition indeed imposes an upper bound on the probability of finding the presence of access to the original work by the model.
 A lower value of $\epsilon$ enhances the effectiveness of this upper bound, and, in the limit of $\epsilon\rightarrow 0$, the evidentiary value of similarity in inferring access vanishes.
@@
 However, the way in which the NAF condition operates in this context is ethically debatable.
-The capacity to generate content with comparable likelihood—regardless of whether access occurred—impedes the determination of whether a work was produced with access based on the generation outcome.
-Intuitively, since a generative model satisfying the NAF condition is likely to produce a work with the same level of similarity even without access to the copyrighted work, the inferential value of similarity in establishing access by the AI model is reduced.
-It is questionable whether such an approach is justified, especially considering the demand for transparency in training data as advocated by researchers (Sag, 2023; Zhang et al., 2024) and reflected in the EU AI Act, [^9] Regulation (EU) 2024/1689 of the European Parliament and of the Council of 13 June 2024 laying down harmonised rules on artificial intelligence and amending Regulations (EC) No 300/2008, (EU) No 167/2013, (EU) No 168/2013, (EU) 2018/858, (EU) 2018/1139 and (EU) 2019/2144 and Directives 2014/90/EU, (EU) 2016/797 and (EU) 2020/1828 (Artificial Intelligence Act), Official Journal of the European Union, L 1689, 12 July 2024, pp.
-1–144.
-which imposes transparency requirements on developers of certain AI systems, mandating that they make available summaries of the training data used.
-Furthermore, since $\Gamma$ is bounded below by the degree of belief $\delta_{k}$ that the model had access to the specific original work in question, given available evidence of access other than similarity, the upper bound on the degree of belief about access becomes increasingly insignificant as stronger evidence of access to the specific original work by the model is presented (e.g., unequivocal evidence of the work’s inclusion in its training data) or as it becomes common practice to include broader sets of copyrighted materials in the training of AI models in general, both of which can lead to higher values of $\delta_{k}$.
-5. Discussion
-Probabilistic analysis has been shown to enhance the understanding of copyright jurisprudence and its interaction with emerging technologies.
-In particular, a mathematical proof demonstrated that the controversial inverse ratio rule holds, provided it is appropriately defined.
-Furthermore, the NAF condition was examined as a mitigation strategy for the copyright infringement risks posed by generative AI, revealing ethical concerns and limitations in its efficacy when addressing the retrospective problem adjudicators face in court, where they must determine whether infringement has occurred based on available evidence.
-There are certain limitations to the analysis presented in this paper.
-While the analysis is based on a probabilistic view of legal fact-finding, this perspective has faced criticisms (Tribe, 1971; Allen and Stein, 2013), although it continues to receive support (Cheng, 2013; Burtis et al., 2017).
-The validity of probabilistic approaches in legal fact-finding in general is widely debated in legal literature and falls beyond the scope of this paper, except for the relation between the results concerning the inverse ratio rule and the conjunction problem.
-Furthermore, the analysis abstracts from certain details involved in AI generation, such as the content of user-provided prompts, which may influence the identification of direct or indirect infringers (Lee et al., 2024) and be important in actual cases.
-Several related directions for future work remain to be explored.
-Although this paper has paid particular attention to the inverse ratio rule—especially compatible with probabilistic reasoning due to its direct connection to the assessment of evidence—the presented approach has broader potential applications within copyright jurisprudence, as many copyright principles can be understood through the lens of evidentiary rules (Lichtman, 2003).
-Furthermore, while the analysis has addressed the NAF condition, future research should focus on developing alternative copyright risk mitigation strategies beyond the NAF condition and conducting a more comprehensive analysis of how generative AI technology interacts with copyright law.
+Intuitively, when the NAF condition is satisfied with small $\epsilon$, training with access to the original work does not increase the model’s probability of generating any particular output much (not more than a multiplicative factor $e^{\epsilon}$) compared to training without access.
+This, in turn, ensures that the model is not significantly more likely to produce outputs similar to the original work relative to a version of the model that did not have access.
+Hence, one cannot strongly infer that the model must have been trained on the original work just because it produces a similar output.
+However, this approach may contradict the demand for transparency in training data, as advocated by researchers (Sag, 2023; Yoo, 2024).
+Furthermore, $\Gamma$ is bounded below by the degree of belief $\delta_{k}$ that the model had access to the specific original work in question, given available evidence of access other than similarity.
+The upper bound on the degree of belief about access becomes less significant as the baseline belief $\delta_{k}$ increases, either through stronger evidence of access to the original work by the model (e.g., unequivocal evidence of its inclusion in the training data) or through the increasingly common practice of training AI models on broader sets of copyrighted materials.
+4. Concluding Remarks
+Probabilistic analysis has been shown to enhance the understanding of copyright jurisprudence and its interaction with generative AI.
+However, there are certain limitations to the analysis presented in this paper.
+While the analysis is based on a probabilistic view of legal fact-finding, this perspective has faced sustained criticism (Tribe, 1971; Allen and Stein, 2013); evaluating the validity of probabilistic approaches in legal fact-finding in general falls beyond the scope of this paper.
+Furthermore, the analysis abstracts from certain details involved in AI generation, such as the content of user-provided prompts, which may influence the identification of direct or indirect infringers (Lee et al., 2024).
+Several future directions remain to be explored.
+Firstly, the presented approach has wider potential applications within copyright jurisprudence, where many legal doctrines can be understood from evidentiary perspectives (Lichtman, 2003).
+In addition, a variety of formal and computational methodologies (Verheij et al., 2015; Halpern, 2017) for evidentiary and probabilistic reasoning beyond the simple framework used here could be applied to further refine and operationalize the probabilistic approach.
+Lastly, future research should analyze and develop copyright risk mitigation strategies alternative to the NAF condition.
 Acknowledgements.
@@
 Appendix A The Inverse Ratio Rule with Strict Inequality
-Proposition 3.2 was proved by assuming only very general relationships among the $\alpha_{i}$’s as well as among the $\beta_{i,j}$’s.
-It does not specify, for example, how the value of $\beta_{i,j}$ changes when one index is increased while the other is decreased simultaneously, nor does it address the relative magnitude of the changes in the values of $\alpha_{i}$ and $\beta_{i,j}$.
-Proposition A.1 demonstrates that, under certain assumptions regarding these details, the inverse ratio rule holds with strict inequality.
-The inequality $\alpha_{\mu^{\prime}}/\alpha_{\mu^{\prime}-1}\leq\beta_{\mu^{\prime}-1,j}/\beta_{\mu^{\prime},j^{\prime}}$ assumed in the first part of Proposition A.1 roughly states that the increase in the degree of belief in access resulting from stronger evidence of access (excluding similarity) is large relative to the changes in the degrees of belief in both access and copying that arise from different levels of similarity.
-On the other hand, the inequality $\beta_{i^{\prime},\kappa^{\prime}}/\beta_{i,\kappa^{\prime}-1}\leq\alpha_{i}/\alpha_{i^{\prime}}$ in the second part of the proposition roughly means that the increase in the degree of belief in copying due to a higher level of similarity is substantial relative to the possible increase (or decrease, in which case the inequality is trivially satisfied) in the degree of belief in access with stronger evidence of access paired with a showing of a lower level of similarity.
-Importantly, these relations are local in the sense that they only have to hold for some specific indices for the proposition to apply, and it is not necessary for them to be universally true across all indices $i$ and $j$.
-It is quite plausible, therefore, that the proposition apply in some actual instances.
+The inverse ratio rule holds with strict inequalities under moderate additional assumptions (Proposition A.1).
 Proposition A.1.
-Assuming that the random variables are defined as in Definition 3.1, the following strong inverse ratio rule holds for any $\lambda\in(0,1)$:
-  • Suppose $j>j^{\prime}$ and $\mu^{\prime}>1$ where $\mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$.
-Then, $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}<\mu^{\prime}$ if $\alpha_{\mu^{\prime}}/\alpha_{\mu^{\prime}-1}\leq\beta_{\mu^{\prime}-1,j}/\beta_{\mu^{\prime},j^{\prime}}$.
-  • Suppose $i>i^{\prime}$ and $\kappa^{\prime}>1$ where $\kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$.
-Then, $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}<\kappa^{\prime}$ if $\beta_{i^{\prime},\kappa^{\prime}}/\beta_{i,\kappa^{\prime}-1}\leq\alpha_{i}/\alpha_{i^{\prime}}$.
-Appendix B Relation to the Conjunction Problem
-The conjunction problem is frequently cited to criticize probabilistic approaches to legal fact-finding (Pardo and Allen, 2008).
-It refers to the issue that when a claim consists of multiple elements that must each be established by evidence, the probability of the conjunction of these elements may not reach a required threshold $\lambda$, even if the probability of each individual element exceeds $\lambda$.
-More concretely, suppose there are two events, $E_{1}$ and $E_{2}$, that need to be established to make a certain claim.
-The probability that both occurs is the product of the probabilities of the two elements, $P(E_{1}\cap E_{2})=P(E_{1}|E_{2})P(E_{2})$, which is smaller than the probability of either individual element.
-Thus, even if both elements have been shown to have probabilities higher than $\lambda$ individually, $P(E_{1}|E_{2}),P(E_{2})>\lambda$, it can be the case that the probability of both requirements being satisfied simultaneously is below the threshold, $P(E_{1}\cap E_{2})<\lambda$.
-A solution offered by the probabilistic camp is to require that the probability of the conjunction, in addition to the individual probability of each element, exceeds the threshold in order to make an affirmative finding (Nance, 1986; Friedman, 1992).
-However, critics object to this solution for its counterintuitiveness (Pardo and Allen, 2008).
-Such objection may appear to affect the results in this paper regarding the inverse ratio rule, which relies on how the probability of copying being a product of two terms: the probability of copying given the presence of access and the probability of access.
-However, this is not the case because the situation here differs from those considered in the context of the conjunction problem.
-Under the Arnstein test, it is not required to prove access to the original work as a separate element from copying.
-Instead, the plaintiff merely needs to show that the probability of copying reaches the threshold, $P(C=1|S=i,\,EA=j)>\lambda$.
-As shown in the proof of Proposition 3.1, due to the logical relationship between copying and access (Definition 3.1 (1)), this probability can be expressed as the product of two terms, the probability of copying and that of access, leading to:
-$\begin{split}&P(C=1|S=i,\,EA=j)>\lambda\\ &\Leftrightarrow P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)>\lambda.\end{split}$
-This product arises merely as a consequence of the logical relationship between copying and access, rather than from adopting the controversial solution that necessitates the probability of the conjunction reaching the threshold, which can be formally described as:
-$P(C=1,\,A=1|S=i,\,EA=j)>\lambda.$
-Therefore, the argument presented here is not directly influenced by discussions surrounding the conjunction problem.
-Appendix C Proofs of Propositions
-C.1. Proof of Proposition 3.2
-Using the logical relationship between copying and access (Definition 3.1 (1)) as well as $EA$’s independence (Definition 3.1 (4)), for any pair $(i,j)$,
-$\begin{split}&P\left(C=1|S=i,\,EA=j\right)\\ &=P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)\\ &\quad+P\left(C=1|A=0,\,S=i\right)P\left(A=0|S=i,\,EA=j\right)\\ &=P\left(C=1|A=1,\,S=i\right)P\left(A=1|S=i,\,EA=j\right)\\ &\quad+0\cdot P\left(A=0|S=i,\,EA=j\right)\\ &=\alpha_{i}\beta_{i,j}.\end{split}$
-The first element of the inverse ratio rule is proved by showing contradiction. Assume $\mu>\mu^{\prime}$ where:
-$\begin{gathered}\mu:=\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\\ \mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}.\end{gathered}$
-Then, by assumptions about the ordering of the terms of the form $\alpha_{i}$ and $\beta_{i,j}$,
-$\begin{split}\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j}\leq\alpha_{\mu}\beta_{\mu,j}\end{split}.$
+The following strong inverse ratio rule holds for any $\lambda\in(0,1)$: Suppose $j>j^{\prime}$ and $\mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}>1$.
+If $\tfrac{\alpha_{\mu^{\prime}}}{\alpha_{\mu^{\prime}-1}}\leq\tfrac{\beta_{\mu^{\prime}-1,j}}{\beta_{\mu^{\prime},j^{\prime}}}$, then $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}<\mu^{\prime}$.
+Similarly, suppose $i>i^{\prime}$ and $\kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}>1$.
+If $\tfrac{\beta_{i^{\prime},\kappa^{\prime}}}{\beta_{i,\kappa^{\prime}-1}}\leq\tfrac{\alpha_{i}}{\alpha_{i^{\prime}}}$, then $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}<\kappa^{\prime}$.
+Appendix B Proofs
+Proof of Proposition 2.2.
+From the logical relationship between copying and access as well as $EA$’s conditional independence, we have $P\left(C=1|S=i,\,EA=j\right)=\alpha_{i}\beta_{i,j}$ for any $(i,j)$.
+The first element of the inverse ratio rule is proved by showing contradiction.
+Assume $\mu>\mu^{\prime}$ where $\mu:=\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}$ and $\mu^{\prime}:=\min\{i:P(C=1|S=i,\,EA=j^{\prime})>\lambda\}$.
+Then, by assumptions about the ordering of the terms of the form $\alpha_{i}$ and $\beta_{i,j}$, the inequalities $\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j}$ hold.
 However, this implies that $\mu^{\prime}$, which is strictly smaller than $\mu$, would have achieved $P(C=1|S=\mu^{\prime},\,EA=j)>\lambda$, contradicting the assumption that $\mu$ was the minimum that achieves this.
-The proof for the second element of the inverse ratio rule is similar. Assume $\kappa>\kappa^{\prime}$ where:
-$\begin{gathered}\kappa:=\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\\ \kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}.\end{gathered}$
-Same as above, the following inequalities hold:
-$\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa}$
-which results in contradiction with the assumption of the minimality of $\kappa$.
-C.2. Proof of Proposition 4.2
-By the NAF condition (Definition 4.1 (1)) and conditional independence (Definition 4.1 (3)):
+The proof for the second element of the inverse ratio rule is similar.
+Assume $\kappa>\kappa^{\prime}$ where $\kappa:=\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}$ and $\kappa^{\prime}:=\min\{j:P(C=1|S=i^{\prime},\,EA=j)>\lambda\}$.
+Then, the inequalities $\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}}$ hold, which is in contradiction with the assumption of minimality of $\kappa$.
+∎
+Proof of Proposition 3.2.
+By the NAF condition and the conditional independence of $EA_{M}$:
 $\begin{split}&P(A_{M}=1|S=i,\,EA_{M}=k)\\ &=\frac{P(S=i|A_{M}=1,\,EA_{M}=k)P(A_{M}=1|EA_{M}=k)}{P(S=i|EA_{M}=k)}\\ &\leq\frac{e^{\epsilon}P(S=i|A_{M}=0,\,EA_{M}=k)P(A_{M}=1|EA_{M}=k)}{P(S=i|EA_{M}=k)}\\ &=\frac{e^{\epsilon}\frac{P(A_{M}=0|S=i,\,EA_{M}=k)P(S=i|EA_{M}=k)}{P(A_{M}=0|EA_{M}=k)}P(A_{M}=1|EA_{M}=k)}{P(S=i|EA_{M}=k)}\\ &=\frac{e^{\epsilon}(1-P(A_{M}=1|S=i,\,EA_{M}=k))P(A_{M}=1|EA_{M}=k)}{1-P(A_{M}=1|EA_{M}=k)}.\end{split}$
-By rearranging terms,
-$\begin{split}P(A_{M}=1|S=i,\,EA_{M}=k)&\leq\frac{e^{\epsilon}P(A_{M}=1|EA_{M}=k)}{1+P(A_{M}=1|EA_{M}=k)(e^{\epsilon}-1)}\\ &=\frac{e^{\epsilon}\delta_{k}}{1+\delta_{k}(e^{\epsilon}-1)}\\ &=\Gamma(\epsilon,\delta_{k}).\end{split}$
-Taking the derivative of $\Gamma(\epsilon,\delta)$ with respect to $\epsilon$ and $\delta$, respectively,
-$\begin{split}&\frac{d\Gamma}{d\mathcal{\epsilon}}=\frac{e^{\epsilon}\delta(1-\delta)}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0\\ &\frac{d\Gamma}{d\delta}=\frac{e^{\epsilon}}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0.\end{split}$
-Therefore, $\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$.
-The inequality $\delta\leq\Gamma(\epsilon,\delta)$ follows from the observation that:
-$\Gamma(\epsilon,\delta)-\delta=\frac{\delta(e^{\epsilon}-1)(1-\delta)}{1+\delta(e^{\epsilon}-1)}\geq 0$
-where the inequality follows from the fact that each term in the numerator and denominator is non-negative because $1\geq\delta\geq 0$ and $e^{\epsilon}>1$.
-Finally, the limit as $\epsilon$ approaches zero can be computed straightforwardly as:
-$\lim_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\frac{1\cdot\delta}{1+\delta\cdot 0}=\delta.$
-C.3. Proof of Proposition A.1
-As in the proof of Proposition 3.2,
-$P(C=1|S=i,\,EA=j)=\alpha_{i}\beta_{i,j},\,\forall(i,j).$
-For the first part, observe:
-$\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}.$
+Rearranging the terms, we obtain the upper bound $\Gamma(\epsilon,\delta_{k})$.
+Taking the derivatives of $\Gamma(\epsilon,\delta)$, we get $\frac{\partial\Gamma}{\partial\mathcal{\epsilon}}=\frac{e^{\epsilon}\delta(1-\delta)}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0$ and $\frac{\partial\Gamma}{\partial\delta}=\frac{e^{\epsilon}}{(1+\delta(e^{\epsilon}-1))^{2}}\geq 0$.
+Thus, $\Gamma(\epsilon,\delta)$ is increasing in $\epsilon$ and $\delta$.
+The inequality $\delta\leq\Gamma(\epsilon,\delta)$ follows from the observation that $\Gamma(\epsilon,\delta)-\delta=\frac{\delta(e^{\epsilon}-1)(1-\delta)}{1+\delta(e^{\epsilon}-1)}\geq 0$, which holds since $1\geq\delta\geq 0$ and $e^{\epsilon}>1$.
+Finally, the limit as $\epsilon$ approaches zero can be computed straightforwardly as: $\lim\limits_{\epsilon\rightarrow 0}\Gamma(\epsilon,\delta)=\frac{1\cdot\delta}{1+\delta\cdot 0}=\delta$.
+∎
+Proof of Proposition A.1.
+As in the proof of Proposition 2.2, $P(C=1|S=i,\,EA=j)=\alpha_{i}\beta_{i,j}$ for any pair $(i,j)$.
+For the first part, observe: $\lambda<\alpha_{\mu^{\prime}}\beta_{\mu^{\prime},j^{\prime}}\leq\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}$.
 In particular, we have $\lambda<\alpha_{\mu^{\prime}-1}\beta_{\mu^{\prime}-1,j}$.
 This implies $\min\{i:P(C=1|S=i,\,EA=j)>\lambda\}\leq\mu^{\prime}-1<\mu^{\prime}$, which completes the proof of the first part.
-Similarly, for the second part, we have:
-$\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}-1}.$
+Similarly, for the second part, we have: $\lambda<\alpha_{i^{\prime}}\beta_{i^{\prime},\kappa^{\prime}}\leq\alpha_{i}\beta_{i,\kappa^{\prime}-1}$.
 This implies $\min\{j:P(C=1|S=i,\,EA=j)>\lambda\}\leq\kappa^{\prime}-1<\kappa^{\prime}$, which completes the proof.
+∎